An individual has $25,000 invested in a stock with a beta of 0.5 and another $80,000 invested in a stock with a beta of 2.0. If these are the only two investments in her portfolio, what is her portfolio's beta? Round your answer to two decimal places.

Assume that the risk-free rate is 7% and the expected return on the market is 12%. What is the required rate of return on a stock with a beta of 2.4? Round your answer to two decimal places.

Assume that the risk-free rate is 5.5% and the market risk premium is 3%.

a. What is the expected return for the overall stock market? Round your answer to two decimal places. %

b. What is the required rate of return on a stock with a beta of 2? Round your answer to two decimal places. %
Answer #1

1.

Portfolio Beta = [25,000(0.50) + 80,000(2.0)]/(105,000)

Portfolio Beta = 1.64

2.

Using CAPM Model,

Required Rate = 0.07 + 2.4(0.12 - 0.07)

Required Rate = 19.0%

3.

Expected Return of Market = 0.055 + 0.03 = 8.3%

Required Rate = 0.055 + 2(0.03) = 11.5%